OpenStudy (anonymous):

Which of the following is a factor of x^2- 21x + 110? (1 point) Question 3 options: 1) x + 11 2) x - 10 3) x - 5 4) None of the above

OpenStudy (mrnood):

when you factorise you can see that the two constant terms must multiply to be the same as the last term in the quadratic so in this case the last term is +110 so if the first option is correct (x+11) then what is the other factor? (x+a) where 11 x a = 110

OpenStudy (mrnood):

so if the second option is correct (x-10) then what is the other factor? (x+a) where -10 x a = +110

OpenStudy (mrnood):

iso if the 3rd option is correct (x-5) then what is the other factor? (x+a) where -5 x a = +110

OpenStudy (mrnood):

in the above I have written x for 'times' (do not confuse it with x the variable) sorry for confusion

OpenStudy (mrnood):

so if the first option is correct (x+11) then what is the other factor? (x+a) where 11 * a = 110 so if the second option is correct (x-10) then what is the other factor? (x+a) where -10 * a = +110 so if the 3rd option is correct (x-5) then what is the other factor? (x+a) where -5 * a = +110

OpenStudy (anonymous):

im not sure whats the answer @MrNood